Find the slope and y-intercept of the line that is perpendicular to y=1/3x-3 and passes through the point (-4, 3).

Answer by mananth(16946)

y = 1/ 3 x + -3
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
m1*m2=-1
The slope of the required line will be -3

m= -3 ,point ( -4 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = 12.00 + b
b= -9
m= -3
The required equation is y =-3x-9
m.ananth@hotmail.ca